# A Closer Look at Model Collapse: From a Generalization-to-Memorization Perspective

### v2

@@
Abstract
The widespread use of diffusion models has led to an abundance of AI-generated data, raising concerns about model collapse—a phenomenon in which recursive iterations of training on synthetic data lead to performance degradation. Prior work primarily characterizes this collapse via variance shrinkage or distribution shift, but these perspectives miss practical manifestations of model collapse. This paper identifies a transition from generalization to memorization during model collapse in diffusion models, where models increasingly replicate training data instead of generating novel content during iterative training on synthetic samples. This transition is directly driven by the declining entropy of the synthetic training data produced in each training cycle, which serves as a clear indicator of model degradation. Motivated by this insight, we propose an entropy-based data selection strategy to mitigate the transition from generalization to memorization and alleviate model collapse. Empirical results show that our approach significantly enhances visual quality and diversity in recursive generation, effectively preventing collapse. The source code is available at https://github.com/shilianghe007/Model_Collapse.git
@@ 1 Introduction
As generative models, such as diffusion models, become widely used for image synthesis and video generation, a large volume of generated data has emerged on the Internet. Since state-of-the-art diffusion models can generate realistic content that even humans cannot easily distinguish, the training datasets for next-generation models will inevitably contain a significant proportion of synthetic data. Figure˜1 illustrates this self-consuming loop, where at each iteration [^1] This paper uses “iteration” to denote a full cycle of training and generation, rather than a gradient update., data generated by the current model is subsequently used to train the new model for the next iteration. Unfortunately, several recent studies have demonstrated that recursively training models on datasets contaminated by AI-generated data leads to performance degradation across these self-consuming iterations—even when synthetic data comprises only a small fraction of the dataset (Dohmatob et al., 2024b). This phenomenon, termed model collapse in prior work, poses a significant threat to the future development and effectiveness of generative models.
@@
As surveyed by Schaeffer et al. (2025), recent studies have identified various collapse behaviors that depend on the performance metrics employed. A series of papers (Shumailov et al., 2024; Kazdan et al., 2024; Bertrand et al., 2024; Suresh et al., 2024) reveal the model collapse phenomenon through the variance of the learned distribution. They empirically and theoretically show that the model continually loses information in the distribution tail, with variance tending towards 0. Another line of work (Dohmatob et al., 2024a; Gerstgrasser et al., 2024; Bertrand et al., 2024; Dohmatob et al., 2024c; Fu et al., 2024) investigates the issue from the perspective of population risk or distribution shifts. These studies observe that the generated distribution progressively deviates from the underlying distribution, causing the model’s population risk to increase throughout the recursive process. Numerous studies (Alemohammad et al., 2024a; Bohacek and Farid, 2023; Zhang et al., 2025b) also report that models begin generating hallucinated data. Despite significant theoretical insights regarding variance dynamics, the reduction of variance to negligible levels typically occurs only after an extremely large number of iterations. As noted by Schaeffer et al. (2025); Kazdan et al. (2024), the collapse progresses at such a slow pace that it is rarely a practical concern in real-world applications. In contrast, the visual quality and diversity of generated images deteriorate rapidly. Furthermore, although population risk or distribution shifts offer a holistic view of performance degradation, they do not adequately characterize specific collapse behaviors.
@@
Accordingly, this paper conducts an in-depth investigation into the collapse dynamics of diffusion models and identifies a generalization-to-memorization transition occurring across successive iterations. Specifically, during early iterations, the model demonstrates a strong capability to generate novel images distinct from those in the training set, but gradually shifts towards memorization in later iterations, merely replicating training images. This transition significantly reduces the diversity of generated content and results in higher FID scores. Moreover, directly reproducing images from training datasets may raise copyright concerns (Ross et al., 2024; Zhang et al., 2024). Furthermore, we empirically reveal a strong linear correlation between the generalizability of the trained model and the entropy of its training dataset. As iterations progress, the entropy of the data distribution sharply decreases, directly resulting in a decline in the model’s generalizability, which illustrates a clear transition from generalization to memorization. Motivated by these empirical findings, we propose entropy-based selection methods to construct training subsets from candidate pools. Extensive experimental validation demonstrates that our proposed methods effectively identify high-entropy subsets, thereby decelerating the generalization-to-memorization transition. Additionally, our approach achieves superior image quality and lower FID scores in recursive training loops compared to baseline methods.
Figure 1: High-level depiction of the self-consuming pipeline. Top: Collapse iteration represents the replace paradigm where models are trained solely on synthetic images generated by the previous diffusion model. Middle: In the mitigated iteration, original real data and previously generated data are added to train the next-generation model. Our proposed selection methods construct a training subset and can further mitigate collapse. Bottom Right: Evolution of the generated images.
@@ 2 Background
In this work, we focus on the image generation task. Let $\mathcal{X}$ be the $d$-dimensional image space, $\mathcal{X}\subseteq\mathbb{R}^{d}$ and let $P_{0}$ be a data distribution over the space $\mathcal{X}$. We use bold letters to denote vectors in $\mathcal{X}$. We assume the original training data $\mathcal{D}_{\mathrm{real}}=\{\bm{x}_{\mathrm{real}}^{(1)},\dots,\bm{x}_{\mathrm{real}}^{(N_{0})}\}$ are generated independently and identically distributed (i.i.d.) according to the underlying distribution $P_{0}$, i.e., $\bm{x}_{\mathrm{real}}^{(i)}\sim P_{0}$.
@@
For a given data distribution, diffusion models do not directly learn the probability density function (pdf) of the distribution; instead, they define a forward process and a reverse process, and learn the score function utilized in the reverse process. Specifically, the forward process (Ho et al., 2020) progressively adds Gaussian noise to the image, and the conditional distribution of the noisy image is given by: $p(\bm{x_{t}}|\bm{x_{0}})=\mathcal{N}\left(\bm{x_{t}};\sqrt{\bar{\alpha}_{t}}\bm{x_{0}},(1-\bar{\alpha}_{t})\bm{I}\right)$, where $\bar{\alpha}_{t}$ is the scale schedule, $\bm{x_{0}}$ is the clean image drawn from $P_{0}$, and $\bm{x_{t}}$ is the noisy image. This forward can also be described as a stochastic differential equation (SDE) (Song and Ermon, 2019): $d\bm{x}=f(\bm{x},t)dt+g(t)d\bm{w}$, where $f(\cdot,t):\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the vector-valued drift coefficient, $g(t)\in\mathbb{R}$ is the diffusion coefficient, and $w$ is a standard Brownian motion. This SDE has a corresponding reverse SDE as $d\bm{x}=[f(\bm{x},t)-g^{2}(t)\nabla_{\bm{x}}\log p_{t}(\bm{x})]dt+g(t)d\bm{w}$, where $dt$ represents a negative infinitesimal time step, driving the process from $t=T$ to $t=0$. The reverse SDE enables us to gradually convert a Gaussian noise to a clean image $\bm{x}\sim P_{0}$.
@@
$\mathbb{E}_{t\sim\mathcal{U}(0,T)}\mathbb{E}_{p_{t}(\bm{x})}\left[\lambda(t)\left\|\nabla_{\bm{x}}\log p_{t}(\bm{x})-\bm{s}_{\theta}(\bm{x},t)\right\|_{2}^{2}\right],$
and can be efficiently optimized with score matching methods such as denoising score matching (Vincent, 2011).
@@
Following the standard setting of model collapse (Alemohammad et al., 2024b; Dohmatob et al., 2024b; Kazdan et al., 2024; Dohmatob et al., 2024d, a; Alemohammad et al., 2024a; Feng et al., 2024), we denote the training dataset at the $n$-th iteration as $\mathcal{D}_{n}$. Let $\mathcal{A}(\cdot)$ denote the training algorithm that takes $\mathcal{D}_{n}$ as input and outputs a diffusion model characterized by the distribution $P_{n}$, i.e., $P_{n}=\mathcal{A}(\mathcal{D}_{n})$. In this work, we train the diffusion model from scratch at each iteration. Subsequently, the diffusion model generates a synthetic dataset of size $N_{n}$, denoted by $\mathcal{G}_{n}\sim P_{n}^{N_{n}}$, which is used in subsequent iterations.
Based on the specific way of constructing training datasets, previous studies (Alemohammad et al., 2024a; Kazdan et al., 2024; Dey and Donoho, 2024) distinguish two distinct iterative paradigms:
  • The replaced training dataset. At each iteration, the training dataset consists solely of synthetic data generated by the previous diffusion model, i.e., $\mathcal{D}_{n}=\mathcal{G}_{n-1}$. Several studies (Gerstgrasser et al., 2024; Kazdan et al., 2024; Dey and Donoho, 2024) refer to this as the “replace” paradigm and have demonstrated that under this setting, the variance collapses to 0 or the population risk diverges to infinity.
  • The accumulated training dataset. A more realistic paradigm (Alemohammad et al., 2024a; Gerstgrasser et al., 2024) is to maintain access to all previous data, thereby including both real images and all synthetic images generated thus far, i.e., $\mathcal{D}_{n}=(\cup_{j=1}^{n-1}\mathcal{G}_{j})\cup\mathcal{D}_{\mathrm{real}}$. However, continuously increasing the training dataset size quickly demands substantial computational resources. A practical compromise is to subsample a fixed-size subset from all candidate images, referred to as the “accumulate-subsample” paradigm in Kazdan et al. (2024). Under certain conditions, prior work (Gerstgrasser et al., 2024; Kazdan et al., 2024) have shown that accumulating real and synthetic data mitigates model degradation, preventing population risk from diverging.
This work focuses on the replace and accumulate-subsample paradigms following prior studies of Shumailov et al. (2024); Suresh et al. (2024); Kazdan et al. (2024). Please refer to the Appendix for additional related work.
@@ 3 Model Collapses from Generalization to Memorization
In this section, we empirically demonstrate the transition from generalization to memorization that occurs over recursive iterations, and investigate the underlying factors driving this transition. All experiments in this section are conducted on the CIFAR-10 dataset (Krizhevsky et al., 2009) using a UNet-based DDPM model (Ho et al., 2020), under the replace paradigm, where each model is trained solely on samples generated by the model from the previous iteration. However, in Appendix˜D, we extend the explorations to other datasets and paradigms, where the conclusion remains valid.
@@
To quantify generalization ability, we adopt the generalization score (Yoon et al., 2023; Alaa et al., 2022; Zhang et al., 2024), defined as the average distance between each generated image and its nearest training image:
$\mathtt{GS}(n)\triangleq\mathtt{Dist}(\mathcal{D}_{n},{\mathcal{G}_{n}})=\frac{1}{|\mathcal{G}_{n}|}\sum_{\bm{x}\in\mathcal{G}_{n}}\min_{z\in\mathcal{D}_{n}}\kappa(\bm{x},\bm{z}),$ (1)
where $\kappa(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ denotes a distance metric between two data points. A higher generalization score $\mathtt{GS}(n)$ indicates that the model generates novel images rather than replicating training samples.
@@
Zhang et al. (2024) measure generalizability as the probability that the similarity between a generated image and its nearest training sample exceeds a threshold. Alaa et al. (2022) assesse memorization via a hypothesis test. While definitions of generalizability vary across studies, they are fundamentally similar, relying on nearest neighbor distances.
@@ 3.1 Finding I: Generalization-to-Memorization Transition
A generalization-to-memorization transition is revealed by experiments showing that the model initially generates novel images but gradually shifts to reproducing training samples in later iterations. We conduct iterative experiments on the CIFAR-10 benchmark (Krizhevsky et al., 2009) to illustrate this transition. Figure˜2 visualizes generated samples alongside their nearest neighbors in the training set. With a relatively large sample size, i.e., 32,76832{,}768 real samples as the starting training dataset, the model tends to generalize first and then memorize. At the early iterations, the diffusion model exhibits strong generalization in early iterations, producing high-quality novel images with little resemblance to training samples. However, its generalization ability deteriorates rapidly over successive iterations, and the model can only copy images from the training dataset after several iterations.
Figure 2: The generalization-to-memorization transition. Left: visualization of the generated images ($\mathcal{G}_{n}$) and their nearest neighbors in the training dataset ($\mathcal{D}_{n}$). As the iteration proceeds, the model can only copy images from the training dataset. Right: quantitative results of the generalization score of models over successive iterations. We use different colors to represent different dataset sizes. A smaller dataset has a larger decaying rate and even falls in the memorization regime at the start (Zhang et al., 2024). We use “iteration” to denote a full cycle of training and generation, rather than a gradient update.
To quantitatively validate the generalization-to-memorization transition, we track the generalization score (GS) introduced in Equation˜1, which measures the similarity between the generated images $\mathcal{G}_{i}$ and the corresponding training images $\mathcal{D}_{i}$ at each iteration. We follow the protocol of Zhang et al. (2024) and construct six nested CIFAR-10 subsets of increasing size: $\lvert\mathcal{D}_{\mathrm{1}}\rvert\in\{1{,}024;2{,}048;4{,}096;8{,}192;16{,}384;32{,}768\}$. These subsets span approximately $3\%$ to $65\%$ of the full training set, providing controlled start points that reflect varying degrees of memorization and generalization. As shown in Figure˜2, GS drops almost exponentially with successive iterations, providing strong empirical evidence for the generalization-to-memorization transition. The decline is noticeably slower for larger training subsets, indicating that larger sample sizes preserve generalization longer and delay the onset of memorization. For the smallest dataset of 1,0241{,}024 images, the model enters the memorization regime from the first iteration and remains there throughout.
@@ 3.2 Finding II: The Entropy of the Training Set Shrinks Sharply over Iterations
We identify entropy as the key evolving factor in the training data that drives the transition from generalization to memorization. Prior work (Yoon et al., 2023) interprets generalization in diffusion models as a failure to memorize the entire training set. Zhang et al. (2024) further show that diffusion models tend to generalize when trained on large datasets (e.g., $2^{14}$ images in CIFAR-10) and to memorize when trained on small ones (e.g., $2^{9}$ images). However, since we fix the size of the training dataset for every iteration, the previous conclusion (Zhang et al., 2024) that a larger dataset leads to generalization cannot fully explain the phenomenon observed in Finding I. We hypothesize that although sample size remains constant, the amount of information it contains decreases over time, making it easier for the model to memorize. Based on this hypothesis, we adopt the differential entropy to quantify the information content and complexity of the continuous image distribution.
Definition 1 (Differential Entropy (Cover, 1999)).
@@
However, the density function $f$ of the image distribution is unknown. We use the following Kozachenko-Leonenko (KL) estimator proposed in a well-known paper (Kozachenko, 1987) to empirically estimate $H(\bm{X})$ from a finite set of i.i.d. samples $\mathcal{D}$} drawn from the distribution $P$:
$\hat{H}_{\gamma}(\mathcal{D})=\psi(|\mathcal{D}|)-\psi(\gamma)+\log c_{d}+\frac{d}{|{\mathcal{D}}|}\sum_{\bm{x}\in{\mathcal{D}}}\log\varepsilon_{\gamma}(\bm{x}),$ (2)
where $\psi:\mathbb{N}\to\mathbb{R}$ is the digamma function, i.e., the derivative of the logarithm of the gamma function; $\gamma$ is any positive integer; $c_{d}$ denotes the volume of the unit ball in the $d$-dimensional space; and $\varepsilon_{\gamma}(\bm{x})=\kappa(\bm{x},\bm{x}_{\gamma})$ represents the $\gamma$-nearest neighbor distance, where $\bm{x}_{\gamma}$ is the $\gamma$-th nearest neighbor of $\bm{x}$ in the set $\mathcal{D}$. Prior work (Kybic and Vnucko, 2012) has shown that the KL estimator is asymptotically unbiased and consistent under broad conditions.
@@ 3.3 Finding III: The Correlation between Entropy and Generalization Score
We verify that the generalization score of the trained model is strongly correlated with the entropy of the training dataset. The similar collapsing trends of entropy and the generalization score motivate a deeper investigation into their relationship. In Figure˜4(a), we present a scatter plot of entropy versus generalization score across different dataset sizes and successive iterations, with the y-axis shown on a logarithmic scale. Notably, the entropy of the training dataset exhibits a significant linear relationship with the logarithm of the generalization score. The Pearson correlation coefficient is $0.91$ with a $p$-value near zero, quantitatively confirming the strength of this correlation. Furthermore, the scatter points corresponding to different dataset sizes are all approximately aligned along a single line, suggesting a dataset size-independent relationship between entropy and generalization. Specifically, training datasets with higher entropy consistently yield better generalization in the trained model. We also observe that larger datasets typically have higher entropy and result in better generalization, aligning with the conclusion in Zhang et al. (2024) regarding the connection between dataset size and generalization. For comparison, Figure˜4(b) shows the scatter plot of variance versus generalization score. The correlation appears substantially weaker than in Figure˜4(a), suggesting that variance in the training dataset may not directly influence the generalization performance of the trained model.
Conclusion.

### v3

@@
A Closer Look at Model Collapse: From a Generalization-to-Memorization Perspective
(December 25, 2025)
Lianghe Shi, Meng Wu, Huijie Zhang, Zekai Zhang, Molei Tao1, Qing Qu University of Michigan    $\cdot$    1Georgia Institute of Technology Joint first author    Corresponding author Abstract The widespread use of diffusion models has led to an abundance of AI-generated data, raising concerns about model collapse—a phenomenon in which recursive iterations of training on synthetic data lead to performance degradation. Prior work primarily characterizes this collapse via variance shrinkage or distribution shift, but these perspectives miss practical manifestations of model collapse. This paper identifies a transition from generalization to memorization during model collapse in diffusion models, where models increasingly replicate training data instead of generating novel content during iterative training on synthetic samples. This transition is directly driven by the declining entropy of the synthetic training data produced in each training cycle, which serves as a clear indicator of model degradation. Motivated by this insight, we propose an entropy-based data selection strategy to mitigate the transition from generalization to memorization and alleviate model collapse. Empirical results show that our approach significantly enhances visual quality and diversity in recursive generation, effectively preventing collapse. Keywords: Model Collapse, Synthetic Data, Memorization, Diffusion Model.  Date: December 25, 2025
Correspondence: lhshi@umich.edu
Resources: Project page   $\cdot$   Code
Figure 1: High-level depiction of the self-consuming pipeline. Top: Collapse iteration represents the replace paradigm where models are trained solely on synthetic images generated by the previous diffusion model. Middle: In the mitigated iteration, original real data and previously generated data are added to train the next-generation model. Our proposed selection methods construct a training subset, further mitigating collapse. Bottom Right: Evolution of the generated images.
@@ 1  Introduction
As generative models, such as diffusion models, become widely used for image synthesis and video generation, a large volume of generated data has emerged on the Internet. Since state-of-the-art diffusion models can generate realistic content that even humans cannot easily distinguish, the training datasets for next-generation models will inevitably contain a significant proportion of synthetic data. Figure˜1 illustrates this self-consuming loop, where at each iteration [^1] This paper uses “iteration” to denote a full cycle of training and generation, rather than a gradient update., data generated by the current model is subsequently used to train the new model for the next iteration. Unfortunately, several recent studies have demonstrated that recursively training models on datasets contaminated by AI-generated data leads to performance degradation across these self-consuming iterations—even when synthetic data comprises only a small fraction of the dataset [dohmatob2024strong]. This phenomenon, termed model collapse in prior work, poses a significant threat to the future development and effectiveness of generative models.
@@
As surveyed by [DBLP:journals/corr/abs-2503-03150], recent studies have identified various collapse behaviors that depend on the performance metrics employed. A series of papers [DBLP:journals/nature/ShumailovSZPAG24, DBLP:journals/corr/abs-2410-16713, DBLP:conf/iclr/BertrandBDJG24, DBLP:journals/corr/abs-2412-17646] reveal the model collapse phenomenon through the variance of the learned distribution. They empirically and theoretically show that the model continually loses information in the distribution tail, with variance tending towards 0. Another line of work [DBLP:conf/nips/DohmatobFK24, DBLP:journals/corr/abs-2404-01413, DBLP:conf/iclr/BertrandBDJG24, DBLP:journals/corr/abs-2410-04840, DBLP:conf/icml/Fu [^0] T24] investigates the issue from the perspective of population risk or distribution shifts. These studies observe that the generated distribution progressively deviates from the underlying distribution, causing the model’s population risk to increase throughout the recursive process. Numerous studies [DBLP:conf/iclr/AlemohammadCLHB24, DBLP:journals/corr/abs-2311-12202, DBLP:journals/corr/abs-2502-09963] also report that models begin generating hallucinated data. Despite significant theoretical insights regarding variance dynamics, the reduction of variance to negligible levels typically occurs only after an extremely large number of iterations. As noted by [DBLP:journals/corr/abs-2503-03150, DBLP:journals/corr/abs-2410-16713], the collapse progresses at such a slow pace that it is rarely a practical concern in real-world applications. In contrast, the visual quality and diversity of generated images deteriorate rapidly. Furthermore, although population risk or distribution shifts offer a holistic view of performance degradation, they do not adequately characterize specific collapse behaviors.
@@
Accordingly, this paper conducts an in-depth investigation into the collapse dynamics of diffusion models and identifies a generalization-to-memorization transition occurring across successive iterations. Specifically, during early iterations, the model demonstrates a strong capability to generate novel images distinct from those in the training set, but gradually shifts towards memorization in later iterations, merely replicating training images. This transition significantly reduces the diversity of generated content and results in higher FID scores. Moreover, directly reproducing images from training datasets may raise copyright concerns [DBLP:journals/corr/abs-2411-00113, pmlr-v235-zhang24cn]. Furthermore, we empirically reveal a strong linear correlation between the generalizability of the trained model and the entropy of its training dataset. As iterations progress, the entropy of the data distribution sharply decreases, directly resulting in a decline in the model’s generalizability, which illustrates a clear transition from generalization to memorization. Motivated by these empirical findings, we propose entropy-based selection methods to construct training subsets from candidate pools. Extensive experimental validation demonstrates that our proposed methods effectively identify high-entropy subsets, thereby decelerating the generalization-to-memorization transition. Additionally, our approach achieves superior image quality and lower FID scores in recursive training loops compared to baseline methods.
@@ 2  Background
[…]
@@
[…]
@@
[…]
and can be efficiently optimized with score matching methods such as denoising score matching [DBLP:journals/neco/Vincent11].
@@
Following the standard setting of model collapse [DBLP:journals/corr/abs-2408-16333, dohmatob2024strong, DBLP:journals/corr/abs-2410-16713, DBLP:conf/icml/DohmatobFYCK24, DBLP:conf/nips/DohmatobFK24, DBLP:conf/iclr/AlemohammadCLHB24, feng2024beyond], we denote the training dataset at the $n$-th iteration as $\mathcal{D}_{n}$. Let $\mathcal{A}(\cdot)$ denote the training algorithm that takes $\mathcal{D}_{n}$ as input and outputs a diffusion model characterized by the distribution $P_{n}$, i.e., $P_{n}=\mathcal{A}(\mathcal{D}_{n})$. In this work, we train the diffusion model from scratch at each iteration. Subsequently, the diffusion model generates a synthetic dataset of size $N_{n}$, denoted by $\mathcal{G}_{n}\sim P_{n}^{N_{n}}$, which is used in subsequent iterations.
Based on the specific way of constructing training datasets, previous studies [DBLP:conf/iclr/AlemohammadCLHB24, DBLP:journals/corr/abs-2410-16713, DBLP:journals/corr/abs-2410-22812] distinguish two distinct iterative paradigms:
  • The replaced training dataset. At each iteration, the training dataset consists solely of synthetic data generated by the previous diffusion model, i.e., $\mathcal{D}_{n}=\mathcal{G}_{n-1}$. Several studies [DBLP:journals/corr/abs-2404-01413, DBLP:journals/corr/abs-2410-16713, DBLP:journals/corr/abs-2410-22812] refer to this as the “replace” paradigm and have demonstrated that under this setting, the variance collapses to 0 or the population risk diverges to infinity.
  • The accumulated training dataset. A more realistic paradigm [DBLP:conf/iclr/AlemohammadCLHB24, DBLP:journals/corr/abs-2404-01413] is to maintain access to all previous data, thereby including both real images and all synthetic images generated thus far, i.e., $\mathcal{D}_{n}=(\cup_{j=1}^{n-1}\mathcal{G}_{j})\cup\mathcal{D}_{\mathrm{real}}$. However, continuously increasing the training dataset size quickly demands substantial computational resources. A practical compromise is to subsample a fixed-size subset from all candidate images, referred to as the “accumulate-subsample” paradigm in [DBLP:journals/corr/abs-2410-16713]. Under certain conditions, prior work [DBLP:journals/corr/abs-2404-01413, DBLP:journals/corr/abs-2410-16713] have shown that accumulating real and synthetic data mitigates model degradation, preventing population risk from diverging.
This work focuses on the replace and accumulate-subsample paradigms following prior studies of [DBLP:journals/nature/ShumailovSZPAG24, DBLP:journals/corr/abs-2412-17646, DBLP:journals/corr/abs-2410-16713]. Please refer to the Appendix for additional related work.
@@ 3  Model Collapses from Generalization to Memorization
In this section, we empirically demonstrate the transition from generalization to memorization that occurs over recursive iterations, and investigate the underlying factors driving this transition. All experiments in this section are conducted on the CIFAR-10 dataset [krizhevsky2009learning] using a UNet-based DDPM model [DBLP:conf/nips/HoJA20], under the replace paradigm, where each model is trained solely on samples generated by the model from the previous iteration. However, in Appendix˜D, we extend the explorations to other datasets and paradigms, where the conclusion remains valid.
@@
To quantify generalization ability, we adopt the generalization score [yoon2023diffusion, DBLP:conf/icml/AlaaBSS22, pmlr-v235-zhang24cn], defined as the average distance between each generated image and its nearest training image:
[…]
@@
[pmlr-v235-zhang24cn] measure generalizability as the probability that the similarity between a generated image and its nearest training sample exceeds a threshold. [DBLP:conf/icml/AlaaBSS22] assesse memorization via a hypothesis test. While definitions of generalizability vary across studies, they are fundamentally similar, relying on nearest neighbor distances.
@@ 3.1  Finding I: Generalization-to-Memorization Transition
A generalization-to-memorization transition is revealed by experiments showing that the model initially generates novel images but gradually shifts to reproducing training samples in later iterations. We conduct iterative experiments on the CIFAR-10 benchmark [krizhevsky2009learning] to illustrate this transition. Figure˜2 visualizes generated samples alongside their nearest neighbors in the training set. With a relatively large sample size, i.e., 32,76832{,}768 real samples as the starting training dataset, the model tends to generalize first and then memorize. At the early iterations, the diffusion model exhibits strong generalization in early iterations, producing high-quality novel images with little resemblance to training samples. However, its generalization ability deteriorates rapidly over successive iterations, and the model can only copy images from the training dataset after several iterations.
Figure 2: The generalization-to-memorization transition. Left: visualization of the generated images ($\mathcal{G}_{n}$) and their nearest neighbors in the training dataset ($\mathcal{D}_{n}$). As the iteration proceeds, the model can only copy images from the training dataset. Right: quantitative results of the generalization score of models over successive iterations. We use different colors to represent different dataset sizes. A smaller dataset has a larger decaying rate and even falls in the memorization regime at the start [pmlr-v235-zhang24cn]. We use “iteration” to denote a full cycle of training and generation, rather than a gradient update.
[…]
@@ 3.2  Finding II: The Entropy of the Training Set Shrinks Sharply over Iterations
We identify entropy as the key evolving factor in the training data that drives the transition from generalization to memorization. Prior work [yoon2023diffusion] interprets generalization in diffusion models as a failure to memorize the entire training set. [pmlr-v235-zhang24cn] further show that diffusion models tend to generalize when trained on large datasets (e.g., $2^{14}$ images in CIFAR-10) and to memorize when trained on small ones (e.g., $2^{9}$ images). However, since we fix the size of the training dataset for every iteration, the previous conclusion [pmlr-v235-zhang24cn] that a larger dataset leads to generalization cannot fully explain the phenomenon observed in Finding I. We hypothesize that although sample size remains constant, the amount of information it contains decreases over time, making it easier for the model to memorize. Based on this hypothesis, we adopt the differential entropy to quantify the information content and complexity of the continuous image distribution.
Definition 3.1 (Differential Entropy [cover1999elements]).
@@
However, the density function $f$ of the image distribution is unknown. We use the following Kozachenko-Leonenko (KL) estimator proposed in a well-known paper [kozachenko1987sample] to empirically estimate $H(\bm{X})$ from a finite set of i.i.d. samples $\mathcal{D}$} drawn from the distribution $P$:
[…]